# Prediction intervals for overdispersed binomial endpoints and their application to historical control data

### v2

@@ Abstract
In toxicology, the validation of the concurrent control by historical control data (HCD) has become requirements. This validation is usually done by historical control limits (HCL) which in practice are often graphically displayed in a Sheward control chart like manner. In many applications, HCL are applied to dichotomous data, e.g. the number of rats with a tumor vs. the number of rats without a tumor (carcinogenicity studies) or the number of cells with a micronucleus out of a total number of cells. Dichotomous HCD may be overdispersed and can be heavily right- (or left-) skewed, which is usually not taken into account in the practical applications of HCL.
To overcome this problem, four different prediction intervals (two frequentist, two Bayesian), that can be applied to such data, are proposed. Comprehensive Monte-Carlo simulations assessing the coverage probabilities of seven different methods for HCL calculation reveal, that frequentist bootstrap calibrated prediction intervals control the type-1-error best. Heuristics traditionally used in control charts (e.g. the limits in Sheward np-charts or the mean ± 2 SD) as well a the historical range fail to control a pre-specified coverage probability.
The application of HCL is demonstrated based on a real life data set containing historical controls from long-term carcinogenicity studies run on behalf of the U.S. National Toxicology Program. The proposed frequentist prediction intervals are publicly available from the R package predint, whereas R code for the computation of the Bayesian prediction intervals is provided via GitHub.
Keywords: Bootstrap-calibration, micro-nucleus-test, long-term carcinogenicity studies, OECD test guideline, Sheward control chart, Bayesian hierarchical modeling
@@ 1 Introduction
For several toxicological or pre-clinical study types it is mandatory to store and report the outcome of the negative and sometimes also positive control groups of the experiments conducted before the current study (see e.g. OECD guidelines 471, 487, 489 or EFSA 2024). In the context of the current study, such historical control data (HCD) can be applied in several ways: One possible application is the inclusion of HCD into the test procedure in order to enhance the power of the current test or to reduce individuals in the concurrent control group (Tarone 1982a, Tarone 1982b, Ryan 1993, Kitsche 2012, Bonapersona et al. 2021, Gurjanov et al. 2021, Gurjanov et al. 2024a, Gurjanov et al. 2024ab). Another possible application is the validation of the concurrent negative (or positive) control group by so called historical control limits (HCL) that aim to define the central x % (usually 95 %) of the underlying distribution of the HCD. Consequently, HCL aim to cover the observation(s) of a current control group with the predefined probability (Kluxen et al. 2021, Dertinger et al. 2023, Menssen 2023). This approach is seen as a quality check for the current control group(s) and it is interpreted as a warning signal, if the current control is not covered by the HCL (Vandenberg et al. 2020): Either as a warning for a false-positive result of the current study (if the concurrent control falls below the lower HCL) or as a warning for a false-negative result (if the concurrent control falls above the upper HCL).
@@
Despite the fact, that the HCL based validation of a current control is required by several test guidelines (OECD 471, OECD 473, OECD 487, OECD 490), most guidelines fail to provide reproducible methodology on how to compute HCL. Hence, many heuristics such as the historical range or the mean $\pm$ 2 SD are used for the practical calculation of HCL, but all heuristics have severe theoretical drawbacks. For a discussion of this topic see Menssen 2023. However, this problem can be overcome by the application of prediction intervals which based on historical observations (the HCD) directly aim for the prediction of one (or more) values of the same data generating process (the current control group). Hence several authors recommend the use of prediction intervals for the application of HCL (EFSA 2024, Menssen and Schaarschmidt 2019, Menssen and Schaarschmidt 2022, Menssen 2023, Kluxen et al. 2021, Dertinger et al. 2023).
In toxicology several study types assess dichotomous endpoints: e.g. Carcinogenicity studies (rats with vs. rats without a tumor), the micro nucleus test (cells with vs. cells without a micronucleus) or the recently developed liquid microplate fluctuation Ames MPFTM test following Spiliotopoulos et al. 2024 (numbers of wells with vs. number of wells without a color change). It seems natural to model such data based on the binomial distribution. But, the clustered nature of the HCD (a certain number of experimental units belong to a certain control group) gives rise for systematic between-study variation. With other words, in many cases the variability of the data exceeds the variance that can be modeled by the simple binomial distribution. This effect is called overdispersion (or extra-binomial variation) and can be commonly found in real life HCD (see supplementary materials section 4) meaning that, possible overdispersion has to be accounted for in the calculation of historical control limits (Menssen and Schaarschmdit 2019, Menssen et al. 2024). Hence, the aim of this paper is to provide reproducible methodology for the calculation of prediction intervals for overdispersed binomial endpoints, that satisfactorily cover the number of affected experimental units in a current control group with the desired probability (e.g. 95 %).
@@
The paper is organized as follows: The next section provides an overview about the modeling of overdispersed binomial endpoints. Section 3 provides an overview about several heuristical methods for the calculation of HCL. Furthermore this section proposes four different methods for the calculation of prediction intervals (two frequentist, two Bayesian). Section 4 reviews the properties of two real life HC data bases, that serve as an inspiration for Monte-Carlo simulations regarding the coverage probabilities of the proposed methodology (section 5). The application of the proposed methods is demonstrated in section 6. The manuscript ends with a discussion (section 7) and some conclusions (section 8).
@@ 2 Models for overdispersed binomial data
Routinely, dichotomous data is modeled based on the binomial distribution, such that
$Y\sim Bin(n,\pi)$ (1)
$E(Y)=n\pi$
$var(Y)=n\pi(1-\pi)$
with $\pi$ as the binomial proportion, and $Y$ number of events out of $n$ experimental units.
However, this approach ignores the clustered structure of the historical control data that is usually comprised of $h=1,\ldots,H$ historical control groups of which each contains $1,\ldots,n_{h}$ experimental units (e.g. rats). This clustering gives rise to possible between-study overdispersion, meaning that the variability of the data exceeds the variability of a simple binomial random variable. This effect can be caused by positive correlations between the experimental units within each cluster (Demetrio et al. 2014, McCullagh and Nelder 1989).
For example, in the context of two-year carcinogenicity studies it is likely, that rats within each control group descent from the same batch of animals which are operated by the same lab personnel, but batches, personnel and other factors are likely to change between different studies. If the tumor rate of some batches is systematically above (or below) average, the number of rats per batch or (historical) control group are no longer independent observations of the same simple binomial distribution as given in eq. 1, because they are positively correlated.
One possible way to model overdispersed binomial data is the beta-binomial distribution in which the binomial proportion for each cluster are derived from a beta distribution
$\pi_{h}\sim beta(a,b)$
@@
$Y_{h}\sim Bin(n_{h},\pi_{h}).$
In this setting $E(\pi_{h})=\pi=a/(a+b)$ and $E(Y_{h})=n_{h}\pi$. The variance is
@@
Note, that in the case in which the number of experimental unis within each control group becomes equal ($n_{h}=n_{h^{\prime}}=n$) also the part of the beta-binomial variance that controls the magnitude of overdispersion per control group ($\big{[}1+(n_{h}-1)\rho\big{]}$ in eq. 2) becomes a constant. Consequently, both models are not in contradiction in this special case (Menssen and Schaarschmidt 2019).
@@ 3 Historical control limits for overdispersed binomial data
The calculation of the historical control limits $[l,u]$ given below is based on the observed number of experimental units with an event $y_{h}$ out of a total number of experimental units $n_{h}$ (e.g. rats with a tumor out of a total number of rats). All historical control limits (HCL) are aimed to cover a further number of experimental units with an event $y^{*}$ out of a further total number of experimental units $n^{*}$ with coverage probability
@@
This means that the historical HCL are aimed to approximate the central $100(1-\alpha)$of the underlying distribution of $y^{*}$. But, overdispersed binomial data can become highly right- or left-skewed, meaning that the skeweness increases the more $\pi$ approaches its boundaries (0 or 1) and/or the lower the sample size $n_{h}$ gets (see fig. 1 and 2 of the supplementary material).
Hence, it is crucial, that the desired HCL account for equal tail probabilities (Hoffmann 2003) in a way that
@@
In this case, the HCL are aimed to converge against the true $\alpha/2$ and $21-\alpha/$ 2 quantiles of the true underlying distribution of $y^{*}$ and hence, control its center (Francq et al. 2019, Menssen 2023).
@@ 3.1 Heuristical historical control limits
Although several authors, dissuade from its use (Greim et al. 2003, Kluxen et al. 2021, EFSA 2024, Menssen 2023), HCL based on the historical range are frequently applied in daily toxicological routine . Since such HCL are given by
@@
they are aimed to cover all observations of the underlying distribution and hence, do not aim for a proper approximation of its central $100(1-\alpha)$%.
Another popular method for the calculation of control limits for dichotomous data are the limits used in Sheward np-charts
@@
with $\bar{\pi}=\frac{\sum_{h}y_{h}}{\sum_{h}n_{h}}$. In practical applications $k$ is usually set to 2 or 3 in order to approximate the central 95.4 % or 99.7 % of the underlying distribution (Montgomery 2019). This HCL are strictly based on the assumption that the observations are independent realizations of the same binomial distribution which can be adequately approximated by a normal distribution with mean $n^{*}\bar{\pi}$ and variance $n^{*}\bar{\pi}(1-\bar{\pi})$. Hence, the Sheward np-chart presumes that the law of large numbers is applicable, but neither allows for overdispersion, nor ensures for equal tail probabilities.
@@
with $\bar{y}=\frac{\sum_{h}y_{h}}{H}$ and $SD=\sqrt{\frac{\sum_{h}(\bar{y}-y_{h})^{2}}{H-1}}$. This type of HCL is based on a simple normal approximation that lacks an explicit assumption about the mean-variance relationship and hence heuristically allows for overdispersion (Menssen et al. 2024). However, it is also based on the assumption that all observations have the same variance, which, for dichotomous data, is only the case, if all control groups (historical, current or future) have the same sample size $n_{h}=n^{*}\quad\forall=1,\ldots,H$ (see eq. 2 and 3). Hence, the application of HCL that are computed by the mean $\pm$ $k$ standard deviations to dichotomous HCD with different sample sizes should strictly be avoided (see section 2 of the supplementary material).
@@ 3.2 Bootstrap calibrated prediction intervals
$\frac{\hat{y}^{*}-Y^{*}}{\sqrt{\widehat{var}(\hat{y}^{*}-Y^{*})}}=\frac{\hat{y}^{*}-Y^{*}}{\sqrt{\widehat{var}(\hat{y}^{*})+\widehat{var}(Y^{*})}}=\frac{n^{*}\hat{\pi}-Y^{*}}{\sqrt{\widehat{var}({n^{*}\hat{\pi})+\widehat{var}(Y^{*})}}}$
is approximately normal. In this notation, $\hat{y}^{*}$ is the expected further observation, $\hat{\pi}$ is the estimate for the binomial proportion estimated from the HCD, $n^{*}$ is the sample size for the predicted observation and the standard error of the prediction is given by
$\widehat{se}(\hat{y}^{*}-Y^{*})=\sqrt{\widehat{var}(\hat{y}^{*})+\widehat{var}(Y^{*})}=\sqrt{\widehat{var}({n^{*}\hat{\pi})+\widehat{var}(Y^{*})}}.$
According to Nelson 1982, the corresponding prediction interval for binomial observations is given as
$[l,u]=n^{*}\hat{\pi}\pm z_{1-\alpha/2}\sqrt{\frac{n^{*2}\hat{\pi}(1-\hat{\pi})}{n}+n^{*}\hat{\pi}(1-\hat{\pi})}.$ (9)
This interval is explicitly based on the assumption that the HCD is an unstructured sample of $n$ experimental units and hence does not allow for overdispersion. Furthermore, it is a symmetrical interval that is based on normal approximation and consequently does not account for equal tail probabilities. Hence the prediction interval given in eq. 9 has to be adapted in a way, that allows its application to clustered HCD which possibly exhibits between-study overdispersion and a skewed distribution. This is done in two steps. Firstly, the asymptotic prediction intervals of Nelson is adapted to overdispersion. In a second step these intervals are adapted to possible skeweness using bootstrap calibration.
If between-study overdispersion is modeled based on the quasi-binomial assumption the prediction standard error becomes $\widehat{se}(\hat{y}^{*}-Y^{*})=\sqrt{\frac{\hat{\phi}n^{*2}\hat{\pi}(1-\hat{\pi})}{\sum_{h}n_{h}}+\hat{\phi}n^{*}\hat{\pi}(1-\hat{\pi})}$ with $\hat{\phi}>1$ as an estimate for the between-study overdispersion. According to Menssen and Schaarschmidt 2019, the corresponding prediction interval is given by
@@
If between-study overdispersion is modeled based on the beta-binomial distribution, the prediction standard error becomes
$\widehat{se}(\hat{y}^{*}-Y^{*})=\sqrt{\Big{[}\frac{n^{*2}\hat{\pi}(1-\hat{\pi})}{\sum_{h}n_{h}}+\frac{\sum_{h}n_{h}-1}{\sum_{h}n_{h}}n^{*2}\hat{\pi}(1-\hat{\pi})\hat{\rho}\Big{]}+n^{*}\hat{\pi}(1-\hat{\pi})[1+(n^{*}-1)\hat{\rho}]}$
with $\hat{\rho}$ as an estimate for the intra-class correlation coefficient. The corresponding prediction interval is given by
$[l,u]=n^{*}\hat{\pi}\pm z_{1-\alpha/2}\sqrt{\Big{[}\frac{n^{*2}\hat{\pi}(1-\hat{\pi})}{\sum_{h}n_{h}}+\frac{\sum_{h}n_{h}-1}{\sum_{h}n_{h}}n^{*2}\hat{\pi}(1-\hat{\pi})\hat{\rho}\Big{]}+n^{*}\hat{\pi}(1-\hat{\pi})[1+(n^{*}-1)\hat{\rho}]}.$ (11)
As mentioned above, overdispersed binomial data can be heavily right- or left-skewed, but the prediction intervals in eq. 10 and 11 are still symmetrical intervals. To overcome this problem, a bootstrap calibration algorithm (see box below), that individually calibrates both interval borders was applied.
@@
    Based on the historical data $\boldsymbol{y}$ find estimates for the model parameters $\hat{\boldsymbol{\theta}}$, with $\hat{\boldsymbol{\theta}}=(\hat{\pi},\hat{\phi})$ in the quasi-binomial case and $\hat{\boldsymbol{\theta}}=(\hat{\pi},\hat{\rho})$ in the beta-binomial case
@@
    Based on $\hat{\boldsymbol{\theta}}$, sample $B$ parametric bootstrap samples $\boldsymbol{y}_{b}$ following the same experimental design as the historical data (for sampling algorithms see section 3 of the supplementary material)
@@
    Draw $B$ further bootstrap samples $y^{*}_{b}$ following the same experimental design as the current observations
@@
    Fit the initial model to $\boldsymbol{y}_{b}$ in order to obtain $\hat{\boldsymbol{\theta}}_{b}$
@@
    Based on $\hat{\boldsymbol{\theta}}_{b}$, calculate $\widehat{var}_{b}(n^{*}\hat{\pi})$ and $\widehat{var}_{b}(Y^{*})$
@@
    Calculate the bootstrapped coverage probability $\hat{\psi}_{l}=\sum_{b}I_{b}\text{ with }I_{b}=1\text{ if }l_{b}\leq y^{*}_{b}\text{ and }I_{b}=0\text{ if }y^{*}_{b}<l_{b}$
@@
    Repeat steps 5-7 for the upper prediction border with $\hat{\psi}_{u}=\sum_{b}I_{b}\text{ with }I_{b}=1\text{ if }y^{*}_{b}\leq u_{b}\text{ and }I_{b}=0\text{ if }u_{b}<y^{*}_{b}$
  10. 10.
    Use the corresponding coefficients $q^{calib}_{l}$ and $q^{calib}_{u}$ for interval calculation

### v3

# Prediction intervals for overdispersed binomial endpoints and their application to toxicological historical control data

@@ Abstract
For toxicology studies, the validation of the concurrent control group by historical control data (HCD) has become requirements. This validation is usually done by historical control limits (HCL), which should cover the observations of the concurrent control with a predefined level of confidence. In many applications, HCL are applied to dichotomous data, e.g. the number of rats with a tumor vs. the number of rats without a tumor (carcinogenicity studies) or the number of cells with a micronucleus out of a total number of cells. Dichotomous HCD may be overdispersed and can be heavily right- (or left-) skewed, which is usually not taken into account in the practical applications of HCL.
To overcome this problem, four different prediction intervals (two frequentist, two Bayesian), that can be applied to such data, are proposed. Based on comprehensive Monte-Carlo simulations, the coverage probabilities of the proposed prediction intervals were compared to heuristical HCL typically used in daily toxicological routine (historical range, limits of the np-chart, mean $\pm$ 2 SD). Our simulations reveal, that frequentist bootstrap calibrated prediction intervals control the type-1-error best, but, also prediction intervals calculated based on Bayesian generalized linear mixed models appear to be practically applicable. Contrary, all heuristics fail to control the type-1-error.
The application of HCL is demonstrated based on a real life data set containing historical controls from long-term carcinogenicity studies run on behalf of the U.S. National Toxicology Program. The proposed frequentist prediction intervals are publicly available from the R package predint, whereas R code for the computation of the two Bayesian prediction intervals is provided via GitHub.
Keywords: Bootstrap-calibration, micro-nucleus-test, long-term carcinogenicity studies, OECD test guideline, Shewhart control chart, Bayesian hierarchical modeling
@@ 1 Introduction
For several toxicological and pre-clinical study types it is mandatory to store and report the outcome of the negative and sometimes also positive control groups of the experiments conducted before the current study (see e.g. OECD guidelines 471, 487, 489 or EFSA 2024). In this context it is required to validate the concurrent control group by so called historical control limits (HCL) which, based on the HCD aim to predict the outcome of the current control group with x % (usually 95 %) of confidence (Kluxen et al. 2021, Dertinger et al. 2023, Menssen 2023).
This approach is seen as a quality check for the current control group(s) and it is interpreted as a warning signal, if the current control is not covered by the HCL (Vandenberg et al. 2020): Either as a warning for a false-positive result of the current study (if the concurrent control falls below the lower HCL) or as a warning for a false-negative result (if the concurrent control falls above the upper HCL). With other words,it is desired that the HCL cover the central 95 % of the possible values that can be obtained from the true underlying distribution.
@@
Despite the fact, that the HCL based validation of a current control is required by several test guidelines (e.g. OECD 471, OECD 473, OECD 487, OECD 490), most guidelines fail to provide reproducible methodology on how to compute HCL. Hence, many heuristics such as the historical range (the min and the max of the HCD) or the mean $\pm$ 2 standard deviations are used for the practical calculation of HCL, but suffer from severe theoretical drawbacks (Menssen 2023). However, this problem can be overcome by the application of prediction intervals, which, based on historical observations (the HCD) directly aim to predict one (or more) value(s) which are derived from the same data generating process (the current control group). Hence several authors recommend the use of prediction intervals for the application of HCL (EFSA 2024, Menssen and Schaarschmidt 2019, Menssen and Schaarschmidt 2022, Menssen 2023, Kluxen et al. 2021, Dertinger et al. 2023).
However, for the proper application of prediction intervals, it is crucial to account for the hierarchical design of the historical control data: Due to the fact that certain experimental units belong to a certain control group (blue panels in fig 1), HCD is comprised of at least one level of hierarchy. Sometimes it might happen, that further levels of hierarchy are present within each control (or treatment) group e.g. that certain animals within a control group belong to a certain cage (shape of the dots in 1). Since each trial is comprised of different living experimental units (cells or animals) standardization between the trials (with respect to the experimental units) is only possible up to a certain extent. Furthermore, lab personnel, housing conditions or animal (or cell) suppliers might change between different historical studies.
Figure 1: Scheme of the hierarchical design of toxicological HCD and the current trial. Blue panels: Control groups of different toxicological studies. Grey panels: Treatment groups of the different toxicological studies, which are usually not taken into account for HCD applications. Black dots: Individual experimental units. Shape of the dots: Possible within group randomization.
The fact that different experimental units belong to different control groups gives rise to systematic between study variation, which usually reflects both: Controllable causes (e.g. differences in handling of experimental units) as well as uncontrollable causes (e.g. the biological variation within the population of the experimental units). This means, that the sheer presence of systematic between study variation is not problematic, but it is rather its magnitude that has to be considered (Deringer et al. 2023). However, this matter needs assay specific discussions by the toxicological community which are off-topic of this manuscript.
It has to be noted that the prediction intervals presented in several standard textbooks and other scientific publications (e.g. Hahn and Meeker 1991, Hahn et al. 2017, Wang 2008, Tian et al. 2022, Nelson 1982) ground on the assumption that the historical sample is unstructured and hence comprised of independent and identically distributed observations, meaning that between-study variability is assumed to be absent. As explained above, this assumption does not make sense in the context of toxicological HCD for which systematic between study variation is a fundamental part of the data generating process, and hence, is regularly reported in real life data. For example, HC data sets that are prone to systematic between study variation can be found in Tarone 1982, Carlus et al. 2013, Menssen et al. 2024, Levy et al. 2019, Tug et al. 2024, Dertinger et al. 2023 (for further details see section 4 of the supplementary material).
In toxicology several study types assess dichotomous endpoints: e.g. Carcinogenicity studies (rats with vs. rats without a tumor), the micro nucleus test (cells with vs. cells without a micronucleus) or the recently developed liquid microplate fluctuation Ames MPFTM test following Spiliotopoulos et al. 2024 (numbers of wells with vs. number of wells without a color change).
It seems natural to model such data based on the binomial distribution. But, as described before, the clustered nature of the HCD gives rise for systematic between-study variation. With other words, if systematic between-study variation is present, the variability of the data will exceed the variance that can be modeled by the simple binomial distribution. This effect is called overdispersion (or extra-binomial variation) and can either be modeled based on the quasi-binomial assumption - which relies on the assumption that the between study overdispersion is constant - or based on the beta-binomial distribution in which the between study overdispersion depends on the number of experimental units within each HC group (see section 2).
In the following, we provide four different methods to compute prediction intervals for overdispersed binomial data: Two frequentist intervals that depend on normal approximation and bootstrap calibration as well as two Bayesian intervals of which one is based on Bayesian hierarchical modeling, whereas the other depends on the Bayesian version of a generalized linear random effects model.
@@
The manuscript is organized as follows: The next section provides an overview about the modeling of overdispersed binomial endpoints. Section 3 provides an overview about the calculation of HCL. Several heuristical HCL that are used in daily toxicological routine are given in section 3.1. The proposed prediction intervals are described in sections 3.2 and 3.3. Section 4 reviews the properties of two real life HC data bases, that serve as an inspiration for Monte-Carlo simulations regarding the coverage probabilities of the proposed methodology (section 5). The application of the proposed methods is demonstrated in section 6. The manuscript ends with a discussion (section 7) and conclusions (section 8).
@@ 2 Models for overdispersed binomial data
2.1 Aggregated data
As described above (see fig. 1), HCD is comprised of $h=1,...,H$ historical control groups but further levels of hierarchy can occur within each control group. For example, it is usual in rodent experiments, that within a certain control group, a predefined number of individuals $n_{hi}$ are housed together in one of $i=1,...,I$ cages. Nevertheless, it frequently happens in practice, that dichotomous HCD is documented aggregated on the level of the control groups (omitting the information about the randomization structure within each control group).
A prominent example for this way of reporting is the Historical Control Data Base of the U.S. National Toxicology Program (NTP 2024). In this data base, the total number of rats (or mice) with a certain tumor is presented on control group level (one value per control group) and the information how many rats were housed per cage is only available as meta-data that does not allow to track which rat (or mouse) belonged to a certain cage (as marked by the different shapes in fig. 1). Hence, the aggregation on control group level, leads to the loss of information on the randomization structure within a control group. Consequently, it is not possible to account for potential within control group overdispersion (caused by systematic differences between cages). To account for this type of reporting, also the prediction intervals presented below are based on aggregated data.
If $n_{hi}$ is the number of experimental units per within control group randomization unit (e.g. cages) and $y_{hi}\in\{0,1,...,n_{hi}\}$ is the observed number of successes among them, aggregation on control group level yields
$y_{h}=\sum_{i=1}^{I}y_{hi}$
experimental units with a finding, out of
$n_{h}=\sum_{i=1}^{I}n_{hi}$
experimental units within the whole control group.
2.2 Modeling between-study overdispersion
One possible way to model overdispersion is the beta-binomial distribution in which the binomial proportion for each cluster is derived from a beta distribution
$\pi_{h}\sim Beta(a,b)$ (1)
@@ 2.2 Modeling between-study overdispersion
$Y_{h}\sim Bin(n_{h},\pi_{h}).$ (2)
In this setting $E(\pi_{h})=\pi=a/(a+b)$, $E(Y_{h})=n_{h}\pi$, $\pi$ is the overall binomial proportion and $n_{h}$ is the a priori known and fixed total number of experimental units per historical control group. The variance of the beta-binomial distributed random variable is
@@
Note, that in the case in which the number of experimental unis within each control group becomes equal ($n_{h}=n_{h^{\prime}}$) also the part of the beta-binomial variance that controls the magnitude of overdispersion per control group ($\big{[}1+(n_{h}-1)\rho\big{]}$ in eq. 3) becomes a constant. Consequently, in this special case, both models are not in contradiction to each other (Menssen and Schaarschmidt 2019) and
$\phi\widehat{=}\big{[}1+(n_{h}-1)\rho\big{]}\text{ if all }n_{h}=n_{h^{\prime}}.$ (5)
@@ 3 Historical control limits for overdispersed binomial data
The calculation of the historical control limits $[l,u]$ given below is based on the observed number of experimental units with an event $y_{h}$ out of a total number of experimental units per historical control group $n_{h}$ (e.g. rats with a tumor out of a total number of rats). All historical control limits (HCL) are aimed to cover a further number of experimental units with an observed number of successes $y^{*}=y_{H+1}$ out of a further a priori known total number of experimental units $n^{*}=n_{H+1}$ with coverage probability
@@
Since HCL are computed to validate the current control group, the desired HCL should aim to cover the central $100(1-\alpha)$of the underlying distribution and its limits should approach the lower $\alpha/2$ and the upper $21-\alpha/$ 2 quantile, respectively (see fig. 1 of Francq et al. 2019). It it crucial to ensure for this behavior also if the underlying distribution is skewed (Hoffman 2003). Consequently, the desired interval has to ensure for equal tail probabilities
@@
This means that the control limits practically cut off the lower and the upper $100(\alpha/2)$of the underlying distribution, regardless if this distribution is symmetrical or skewed. For overdispersed binomial data, possible skewness depends on three factors:
  1. 1.
    The skewness increases, the closer $\pi$ approaches its boundaries (0 or 1).
  2. 2.
    The skewness tends to increase with increasing between-study overdispersion (depending on the value for $\pi$).
  3. 3.
    The skewness also tends to increases with decreasing cluster size $n^{*}$ (depending on the values for $\pi$ and the amount of between-study overdispersion).
This behavior is demonstrated in fig. 1 and 2 of the supplementary material.
@@ 3.1 Heuristical historical control limits
Although several authors, dissuade from its use (Greim et al. 2003, Kluxen et al. 2021, EFSA 2024, Menssen 2023, Menssen et al. 2024), HCL based on the historical range are frequently applied in daily toxicological routine (Zarn et al. 2024, WHO 2023, NTP 2018, EFSA 2010). Since such HCL are given by
@@
they aim to cover all realizations from the underlying distribution and hence, do not aim for a proper approximation of its central $100(1-\alpha)$(but can be rather interpreted as a heuristical 100% prediction interval which does not provide practically useful control limits). Note that the min and max of the observed data provide very poor estimates for the true min and max of the underlying distribution, especially if the number of historical observations is low. This is because values in the center of the distribution have a higher chance to occur, whereas the occurrence of more extreme values in the tails of the distribution are less likely. Furthermore it has to be noted, that this interval is explicitly based on the assumption of constant cluster size $n_{h}=n^{*}\text{ }\forall\text{ }h=1,\ldots H$, because the interpretation of the number of observed events makes only sense in relation to the given cluster size (meaning that 5 out of 50 rats with a tumor have a fundamental different interpretation than 5 out of 100 rats). Hence, a sensible interpretation of the historical range is not possible, if the cluster size differs between the studies.
Another popular method for the calculation of control limits for dichotomous data are the limits used in Shewhart np-charts
@@
with $\bar{\pi}=\frac{\sum_{h}y_{h}}{\sum_{h}n_{h}}$. In practical applications $k$ is usually set to 2 or 3 in order to approximate the central 95.4 % or 99.7 % of the underlying distribution (Montgomery 2019). This HCL are strictly based on the assumption that the observations are independent realizations of the same binomial distribution which can be adequately approximated by a normal distribution with mean $n^{*}\bar{\pi}$ and variance $n^{*}\bar{\pi}(1-\bar{\pi})$. Hence, the Shewhart np-chart presumes that the law of large numbers is applicable, ignores the variability in $\bar{\pi}$, fails to account for overdispersion and fails to ensure for equal tail probabilities.
@@
with $\bar{y}=\frac{\sum_{h}y_{h}}{H}$ and $SD=\sqrt{\frac{\sum_{h}(\bar{y}-y_{h})^{2}}{H-1}}$. This type of HCL is based on a simple normal approximation that lacks an explicit assumption about the mean-variance relationship and hence heuristically allows for overdispersion (Menssen et al. 2024). Similar to HCL that are based on the historical range, the mean $\pm$ $k$ standard deviations are explicitly based on the assumption of constant cluster size $n_{h}=n^{*}\quad\forall\quad h=1,\ldots,H$, because for dichotomous endpoints, the number of observed experimental units with an event is only interpretable in the context of the total number of experimental units in the particular control group. Hence, the application of HCL that are computed by the mean $\pm$ $k$ standard deviations to dichotomous HCD with different sample sizes should strictly be avoided (see section 2 of the supplementary material).
@@ 3.2 Bootstrap calibrated prediction intervals
$\frac{n^{*}\hat{\pi}-Y^{*}}{\sqrt{var(n^{*}\hat{\pi})+var(Y^{*})}}\stackrel{{\scriptstyle\text{approx.}}}{{\sim}}N(0,1).$
Here, $Y^{*}$ is the future random variable (of which $y^{*}$ is a realization), $\hat{\pi}$ is the binomial proportion estimated from the HCD and $n^{*}$ is the a priori known number of experimental units of the future control group and $n^{*}\hat{\pi}$ and $Y^{*}$ are assumed to be independent random variables.
If between-study overdispersion is modeled based on the quasi-binomial assumption, the corresponding prediction interval is given by
@@
with $\hat{\phi}>1$ as an estimate for the between-study overdispersion.
If between-study overdispersion is modeled based on the beta-binomial distribution, the corresponding prediction interval is given by
$[l,u]=n^{*}\hat{\pi}\pm z_{1-\alpha/2}\sqrt{\Big{[}\frac{n^{*2}\hat{\pi}(1-\hat{\pi})}{\sum_{h}n_{h}}+\frac{\sum_{h}n_{h}-1}{\sum_{h}n_{h}}n^{*2}\hat{\pi}(1-\hat{\pi})\hat{\rho}\Big{]}+n^{*}\hat{\pi}(1-\hat{\pi})[1+(n^{*}-1)\hat{\rho}]}$ (12)
with $\hat{\rho}$ as an estimate for the intra-class correlation coefficient. Details on methodology for the estimation of $\hat{\pi}$, $\hat{\phi}$ and $\hat{\rho}$ are given in section 3.4).
As mentioned above, overdispersed binomial data can be heavily right- or left-skewed, but the prediction intervals in eq. 11 and 12 are still symmetrical. Furthermore, they depend on normal approximation and hence, tend to have coverage probabilities below the nominal level, especially if the number of historical studies is low.
To overcome this problem, a bootstrap calibration algorithm (see box below), that individually calibrates both interval borders was applied. Note, that the idea behind this algorithm is similar to a Monte-Carlo simulation regarding the prediction intervals coverage probability. Therefore, the bootstrap samples $\boldsymbol{y}_{b}$ that mimic the HCD, as well the bootstrapped future observation $y^{*}_{b}$ are derived from the same data generating process (which depends on the estimates for the model-parameters estimated from the HCD). The algorithm aims to provide substitutes $q_{l}$ and $q_{u}$ for the standard normal quantile used in eq. 11 and 12, such that possible skewness of the underling distribution is taken into account to ensure for equal tail probabilities and enhance the intervals coverage probability.
@@
    Based on the historical data $\boldsymbol{y}=(y_{1},\ldots,y_{H})$ find estimates for the model parameters ($\hat{\pi}$ and $\hat{\phi}$ in the quasi-binomial case and $\hat{\pi}$ and $\hat{\rho}$ in the beta-binomial case
@@
    Based on these estimates, sample $B$ parametric bootstrap samples $\boldsymbol{y}_{b}$ following the same experimental design as the historical data (for sampling algorithms see section 3 of the supplementary material)
@@
    Draw $B$ further bootstrap samples $y^{*}_{b}$ that mimic the possible numbers of success in a future control group of size $n^{*}$ (using the same sampling algorithms as in step 2)
@@
    Based on $\boldsymbol{y}_{b}$ obtain bootstrapped estimates for the model parameters (either $\hat{\pi}_{b}$ and $\hat{\phi}_{b}$ or $\hat{\pi}_{b}$ and $\hat{\rho}_{b}$) and calculate $\widehat{var}_{b}(n^{*}\hat{\pi})$ and $\widehat{var}_{b}(Y^{*})$
@@
    Calculate the bootstrapped coverage probability $\hat{\psi}_{l}=\frac{\sum_{b}}{B}I_{b}\text{ with }I_{b}=1\text{ if }l_{b}\leq y^{*}_{b}\text{ and }I_{b}=0\text{ if }y^{*}_{b}<l_{b}$
@@
    Repeat steps 5-7 for the upper prediction border with $\hat{\psi}_{u}=\frac{\sum_{b}}{B}I_{b}\text{ with }I_{b}=1\text{ if }y^{*}_{b}\leq u_{b}\text{ and }I_{b}=0\text{ if }u_{b}<y^{*}_{b}$
@@
    Use the corresponding coefficients $q^{calib}_{l}$ and $q^{calib}_{u}$ and the estimates obtained from the initial HCD (either $\hat{\pi}$ and $\hat{\phi}$ or $\hat{\pi}$ and $\hat{\rho}$) for interval calculation